# Stellar cores live long and prosper in cuspy dark matter halos

### v1

@@ 4 Stellar density inference
We note that a crucial assertion when applying the Eddington inversion method to rule out cuspy DM halos is the presence of stellar cores, and more specifically, the presence of non-negative inner density slopes. In almeida_stellar_2024, this assertion is supported by fitting a $5^{\mathrm{th}}$-order polynomial to the combined projected surface density profiles of six UFDs from richstein2024. Here, we test the ability to distinguish different inner density slopes with our simulated stellar Plummer profiles in the fiducial setup ($M_{\mathrm{ext}}=M_{\mathrm{IC}}$).
We fit at times $t=(0,1,2,5,10,20,50,100)\,\mathrm{Gyr}$ the full 3D information of the stellar distribution with the $(\alpha,\beta,\gamma)$ double power-law profile (merritt2006; dicintio2014):
@@
Here $\rho_{\mathrm{S}}$ is a characteristic density and $r_{\mathrm{S}}$ a characteristic scale length. The inner and outer regions have logarithmic slopes of $\gamma$ and $\beta$, respectively, with the transition between these regions governed by $\alpha$. A Plummer sphere (Equation 3) has $(\alpha,\beta,\gamma)=(2,5,0)$, whereas for an NFW profile (Equation 1), the parameterization is $(1,3,1)$.
@@
To infer the analytic stellar density profile from the simulation data, we fit the $(\alpha,\beta,\gamma)$ profile to the stellar density, using the same binning as discussed in section 3.
@@
We use hierarchical Bayesian inference to jointly fit the density profiles of all 10 fiducial runs at the aforementioned times: a detailed methodology is provided in Appendix B. This way, we avoid many of the problems — such as overfitting, loss of data heterogeneity, and diminished precision of parameter estimates for small sample sizes — that are commonly associated with inference performed when completely stacking the data into a single set or performing inference on related datasets independently (gelman2006).
@@
An important finding of our analysis is the inability to constrain $\gamma$ as strictly-positive at any time, indicating that despite a priori knowledge that the derivative of the stellar density profile from which our samples are drawn is always $\mathrm{d}\rho/\mathrm{d}r\leq 0$, this cannot be unequivocally said from the measured density profile. By maximizing the data available in the inference, we constrain the population mean and standard deviation of the inner density slope at $t=0\,\mathrm{Gyr}$ as $\mu_{\gamma}=-0.005_{-0.208}^{+0.178}$ and $\sigma_{\gamma}=0.011_{-0.010}^{+0.026}$, respectively, where the uncertainties refer to the 90% confidence interval of the parameter. The time evolution of the full distribution of $\gamma$ is shown in the top-left panel of Figure 4. At all times, $\gamma$ is consistent with a value of 0, however the spread in the distributions means that $\gamma<0$ could be easily inferred: we find that the probability $P(\gamma\leq 0)=0.52$.
@@
For comparison, we fit a non-hierarchical model independently to the 10 fiducial simulations at $t=0\,\mathrm{Gyr}$ (i.e., we fit for the parameters in the $(\alpha,\beta,\gamma)$ profile directly). Unsurprisingly, the constraints on $\gamma$ are significantly weaker when inferring the density model parameters independently for each simulation compared to the partial-stacking allowed for by the hierarchical model, as shown in the top-right panel of Figure 4. While all individual realizations are drawn from the same underlying Plummer distribution with $\gamma=0$, due to sample variance, the probability distribution function (PDF) inferred from each realization differs. In particular, in every realization, the PDF extends to both positive and negative values of $\gamma$.
@@
To demonstrate degeneracies between the parameters of the fitted $(\alpha,\beta,\gamma)$ profile, in the bottom panels of Figure 4 we show contours of the 68% and 95% confidence intervals of the inferred parameters for the hierarchical model. We see that their true values ($\alpha=2,\,\beta=5,\,\gamma=0,\,r_{\mathrm{S}}=0.2\,\mathrm{kpc},\,\mathrm{and}\,\rho_{\mathrm{S}}\simeq 1.5\times 10^{5}\,\mathrm{M}_{\sun}\mathrm{kpc}^{-3}$) are recovered very well in this case. However, $\gamma$ is strongly correlated with the other parameters, especially with $\rho_{\mathrm{S}}$ but also clearly with $\alpha$ and $\beta$.
5 Discussion
@@
When we discuss cores, we use the widely adopted definition of density profiles that are asymptotically flat, i.e., whose inner slope $\mathrm{d}\rho/\mathrm{d}r\rightarrow 0$ as $r\rightarrow 0$. That is, we consider a density profile such as the Plummer profile (Equation 3), or the $(\alpha,\beta,\gamma)$ profile (Equation 7) with $\gamma=0$, to be cored, even though $\mathrm{d}\rho/\mathrm{d}r<0$ for any finite $r$. Applying the Eddington inversion method to rule out cuspy potentials may require a more limited definition ($\mathrm{d}\rho/\mathrm{d}r>0$ at some finite radius) by which Plummer profiles would not be considered cored. However, we have shown that for typical UDFs, measurements of the true inner slope from small numbers of stars are ambiguous.
@@
A known constraint of the Eddington inversion method (almeida_EIM_2024) is that perfect stellar cores with $\gamma\leq 0$ are incompatible with NFW potentials when the velocity distribution is isotropic or radially biased ($\beta_{\mathrm{ani}}\geq 0$). The incompatibility arises from the method producing a DF that is not everywhere non-negative (i.e., there exists an $\mathcal{E}$ such that $f(\mathcal{E})<0$), and is hence unphysical.
@@
If we consider our ICs, which are by definition ergodic and spherically-symmetric, the DF is solely a function of the relative energy (as in Equation 4). Effectively, $f$ tells us the probability of finding a stellar particle in the system with an energy in the infinitesimal volume $\mathrm{d}\mathcal{E}$ about $\mathcal{E}$. In practice however, we are unable to numerically sample from such an infinitesimal volume, and rather sample from a coarsened DF $\bar{f}(\mathcal{E})$ (binney2008) that defines a finite volume of width $\Delta\mathcal{E}$, where $\bar{f}$ is the expectation value, or mean, of $f$ within the volume $\Delta\mathcal{E}$. Critically, whilst $\bar{f}(\mathcal{E})$ may be everywhere non-negative in the volume $\Delta\mathcal{E}$, $f(\mathcal{E})$ in a subregion $\mathrm{d}\mathcal{E}\in\Delta\mathcal{E}$ may not be. The task then becomes ensuring that our approximation
@@
is a good one over the domain $\Omega$.
@@
In creating our ICs, we calculate $f$ at $2\times 10^{3}$ equally-spaced values of $\mathcal{E}$: this defines our coarsened DF [^1] Note that we use interpolation so as to evaluate our DF at any $\mathcal{E}$, and not just at the values used to construct the DF.. We have tested the effect of increasing the number of $\mathcal{E}$ values $f$ is evaluated at by a factor of 10, and performed a two-sample Kolmogorov–Smirnov (KS) test with a two-tailed null hypothesis to assess if there is a meaningful difference between evaluations of the DF. We find that the test statistic is $t_{\mathrm{KS}}=9.5\times 10^{-4}$ and the $p$-value is $\sim 1.0$: the DFs are indistinguishable. Consequently, we are confident that our constructed $\bar{f}$ is a good approximation to the true $f$. This is further supported by the constructed ICs agreeing well with the desired Plummer profile (shown in Figure 4). While negative DFs are clearly unphysical, analytic DFs are only descriptive, which real and finite systems are not required to fully sample.
6 Conclusions
@@
The existence of cored rather than cuspy DM profiles in UFDs would have profound implications, effectively falsifying the standard cosmological model. However, our results challenge previous assertions of evidence for DM cores based on cored stellar density profiles. Firstly, we have shown that cored stellar systems can persist within cuspy halos, provided that they have formed in initial equilibrium, as would be expected for a stellar system forming within an existing DM halo in the standard galaxy formation paradigm.
@@
Secondly, we have demonstrated the limitations of determining the crucial inner slope of the stellar density profiles of UFDs. Even with a conservative sample size and in the absence of observational errors, it is exceptionally difficult to distinguish in UFDs a mildly positive, negative, or zero inner density gradient. This is heightened by the small collection of stars that constitute such a dwarf galaxy, increasing the observed variance in stellar mass density profiles. While future observations may yet lead to a viable application of the Eddington inversion method, we conclude that based on current observations, the existence of stellar cores in UFDs provides no evidence against cuspy DM halos.
@@
J.H. acknowledges the support of the Magnus Ehrnrooth Foundation. A.R. acknowledges the support of the University of Helsinki Research Foundation. J.H. and T.S. acknowledge support by Research Council of Finland grants 354905 and 339127. J.H., A.R., and T.S. acknowledge support by ERC Consolidator Grant KETJU (no. 818930). M.G.W. acknowledges support from the National Science Foundation (NSF) grant AST-2206046. This work used facilities hosted by the CSC – IT Centre for Science, Finland. {contribution}JH: data curation, formal analysis, investigation, methodology, visualization, writing – original draft. AR: formal analysis, investigation, methodology, software, writing – original draft. TS: conceptualization, supervision, writing – original draft. MGW: conceptualization, writing – original draft.
@@ Appendix A Binning of stellar particles
We test the effect of using different numbers of bins when constructing radial profiles on our conclusions, in particular for the inference of the $(\alpha,\beta,\gamma)$ double power-law profile. In all cases, we use bins of equal particle counts (as opposed to equal bin widths) to ensure sufficient and comparable sample sizes within each bin from which we determine statistics, such as the mean density within a bin. In the following, we discuss signal-to-noise (S/N) in the Poisson sense, i.e., $\mathrm{S/N}=N/\sqrt{N}=\sqrt{N}$.
We first note that determining an optimal binning strategy is a result of two competing factors (e.g., tarumi2021): increasing S/N, achieved with increasing the number of particles per bin, and increasing spatial resolution, achieved with increasing the total number of bins. Motivated by observational campaigns of UFDs (e.g., drlica-wagner2015; tan2025), we require that the minimum S/N in each radial bin is $\mathrm{S/N}\simeq 10\implies N\simeq 100$. We thus try binning the inner 95% of stellar particles using 5, 10, 25, and 50 bins of equal particle counts [^2] We also tested a case using 95 bins, however due to the low S/N the profile fits were unable to be recovered.. Already we note that using 5 bins, whilst achieving a high S/N of $\sim 30$, barely resolves the inner regions of the stellar profile $r<r_{\mathrm{S}}$, thus setting a lower limit to the number of bins. Using 50 bins achieves our desired $\mathrm{S/N}\simeq 10$, thus setting our upper limit.
@@
A smaller $d_{\mathrm{M}}$ is indicative of a better recovery of the a priori truth than a higher $d_{\mathrm{M}}$. We find that as a function of bin count, $d_{\mathrm{M}}$ decreases exponentially as $d_{\mathrm{M}}\propto N_{\mathrm{bin}}^{-0.43}$: from $N_{\mathrm{bin}}=25$ to $N_{\mathrm{bin}}=50$, the difference in distance measures is only $\Delta d_{\mathrm{M}}=0.24$, indicating the two strategies recover the a priori truth comparatively well. We thus adopt the binning that provides the higher spatial resolution, namely $N_{\mathrm{bin}}=50$, as the chosen binning strategy in this work.
@@ Appendix B Fitting density profiles
We fit the $(\alpha,\beta,\gamma)$ double power-law profile using hierarchical Bayesian inference implemented in Stan (standevelopmentteam2018), with four chains of 2000 posterior draws each (i.e., post burn-in period). Recognizing that the 10 fiducial simulations describe the same physical system, albeit with different realizations of the particle sampling, we can jointly model the individual density profile parameters, denoted $\bm{\theta}\equiv\left[\log_{10}\rho_{\mathrm{S}},\log_{10}r_{\mathrm{S}},\log_{10}\alpha,\beta,\gamma\right]$, as realizations of a population of possible density profiles that satisfy the exchangeability criterion (gelman2015). The population distribution of density profiles is governed by hyperparameters on each of the latent (or observed) parameters in $\bm{\theta}$ that themselves have weakly-informative priors. The prior distributions for the hyperparameters are given in Table 1, and a graphical representation of the Bayesian model in the form of a directed acyclic graph is given in Figure 5. Each latent parameter in $\bm{\theta}$ we assign a normal distribution, with mean $\mu_{X}$ and standard deviation $\sigma_{X}$ for the respective variable $X\in\bm{\theta}$. To capture any covariance between latent parameters, we directly model the correlation matrix $\Omega=L_{\mathrm{c}}L_{\mathrm{c}}^{\top}$, where $L_{\mathrm{c}}$ is the Cholesky factor [^3] The Cholesky factor of a positive-definite matrix is a lower-triangular matrix $L$ such that $\Sigma=LL^{\top}$. of the correlation matrix. We assign a lewandowski2009 prior on $L_{\mathrm{c}}$ as $L_{\mathrm{c}}\sim\mathrm{LKJCholesky}(\eta)$, with the prior on the correlation matrix thus becoming $p(\Omega|\eta)\propto\det(\Omega)^{\eta-1}$. The variable $\eta$ controls the concentration of the prior about the identity matrix: by setting $\eta=2$ (Table 1), we mildly-favor weak correlations while allowing moderate correlation structure.
@@
After the posterior sampling, we perform standard convergence and posterior-predictive checks to ensure the fits are robust by enforcing a diagnostic value $\hat{R}$ (comparison of between-chain and within-chain variance) of less than 1.05, a high effective sample, and that the number of divergent transitions is less than 1 percent (for details, see vehtari2021; standevelopmentteam2018). We perform an additional prior sensitivity analysis using the power-scaling method described in kallioinen2024, ensuring that the sensitivity diagnostic is less than 0.05 for the latent parameters $\bm{\theta}$ of the model. In doing so, we quantitatively ensure that the obtained posterior distributions are not strongly-dependent on the assumed prior distributions.
As our model takes a number of free parameters, degeneracies between parameters are possible, thus highlighting the importance of using a Bayesian method to propagate parameter uncertainty. Typically, we find that the inner slope $\gamma$ is strongly negatively-correlated with the characteristic density $\log_{10}\rho_{\mathrm{S}}$, as well as being correlated to a lesser degree with both $\alpha$ (positively) and $\beta$ (negatively). Consequently, $\log_{10}\rho_{\mathrm{S}}$ tends to be negatively-correlated with $\alpha$ and positively-correlated with $\beta$.

### v2

@@ IV Stellar density inference
We note that a crucial assertion when applying the Eddington inversion method to rule out cuspy DM halos is the presence of stellar cores, and more specifically, the presence of nonnegative inner density slopes. In Sánchez Almeida et al. (2024b), this assertion is supported by fitting the combined projected surface density profiles of six UFDs from Richstein et al. (2024) and inferring the corresponding phase-space DF. The profiles for all six UFDs are assumed to have the same polynomial shape, which is found to have an inner plateau with a near-zero slope. In an NFW potential, the best-fitting DF producing these profiles is found to contain negative parts, leading the authors to conclude such potential to be inconsistent with the observations. Here, we test the ability to distinguish different inner density slopes with our simulated stellar Plummer profiles in the fiducial setup ($M_{\mathrm{ext}}=M_{\mathrm{IC}}$).
We fit at times $t=(0,1,2,5,10,20,50,100)\,\mathrm{Gyr}$ the full 3D information of the stellar distribution with the $(\alpha,\beta,\gamma)$ double-power-law profile (Merritt et al., 2006; Di Cintio et al., 2014):
@@
Here $\rho_{\mathrm{S}}$ is a characteristic density, and $r_{\mathrm{S}}$ is a characteristic scale length. The inner and outer regions have logarithmic slopes of $\gamma$ and $\beta$, respectively, with the transition between these regions governed by $\alpha$. A Plummer sphere (Equation 3) has $(\alpha,\beta,\gamma)=(2,5,0)$, whereas for an NFW profile (Equation 1), the parameterization is $(1,3,1)$. We fit for all free parameters in Equation 7 ($\alpha,\beta,\gamma,r_{\mathrm{S}},\mathrm{and}\,\rho_{\mathrm{S}}$).
@@
To infer the analytic stellar density profile from the simulation data, we fit the double-power-law profile to the stellar density, using the same binning as discussed in section III.
@@
We use hierarchical Bayesian inference to jointly fit the density profiles of all 10 fiducial runs at the aforementioned times: a detailed methodology is provided in Appendix B. This way, we avoid many of the problems—such as overfitting, loss of data heterogeneity, and diminished precision of parameter estimates for small sample sizes—that are commonly associated with inference performed when completely stacking the data into a single set or performing inference on related datasets independently (Gelman and Hill, 2006).
@@
An important finding of our analysis is the inability to constrain $\gamma$ as strictly positive at any time, indicating that despite a priori knowledge that the derivative of the stellar density profile from which our samples are drawn is always $\mathrm{d}\rho/\mathrm{d}r\leq 0$, this cannot be unequivocally said from the measured density profile. By maximizing the data available in the inference, we constrain the population mean and standard deviation of the inner density slope at $t=0\,\mathrm{Gyr}$ as $\mu_{\gamma}=-0.005_{-0.208}^{+0.178}$ and $\sigma_{\gamma}=0.011_{-0.010}^{+0.026}$, respectively, where the uncertainties refer to the 90% confidence interval of the parameter. The time evolution of the full distribution of $\gamma$ is shown in the top-left panel of Figure 4. At all times, $\gamma$ is consistent with a value of 0; however the spread in the distributions means that $\gamma<0$ could be easily inferred: we find that the probability $P(\gamma\leq 0)=0.52$.
@@
For comparison, we fit a nonhierarchical model independently to the 10 fiducial simulations at $t=0\,\mathrm{Gyr}$ (i.e., we fit for the free parameters in the double-power-law profile directly). Unsurprisingly, the constraints on $\gamma$ are significantly weaker when inferring the density model parameters independently for each simulation compared to the partial stacking allowed for by the hierarchical model, as shown in the top-right panel of Figure 4. While all individual realizations are drawn from the same underlying Plummer distribution with $\gamma=0$, due to sample variance, the probability distribution function (PDF) inferred from each realization differs. In particular, in every realization, the PDF extends to both positive and negative values of $\gamma$.
@@
To demonstrate degeneracies between the parameters of the fitted double-power-law profile, in the bottom panels of Figure 4, we show contours of the 68% and 95% confidence intervals of the inferred parameters for the hierarchical model. We see that their true values ($\alpha=2,\,\beta=5,\,\gamma=0,\,r_{\mathrm{S}}=0.2\,\mathrm{kpc},\,\mathrm{and}\,\rho_{\mathrm{S}}\simeq 1.5\times 10^{5}\,\mathrm{M}_{\sun}\mathrm{kpc}^{-3}$) are recovered very well in this case. However, $\gamma$ is strongly correlated with the other parameters, especially with $\rho_{\mathrm{S}}$ but also clearly with $\alpha$ and $\beta$. This further shows the difficulty in constraining the inner density slope precisely: for instance, even a slight increase in $\rho_{\mathrm{S}}$ favors a more positive inner slope ($\gamma<0$), and vice versa. In other words, the degeneracy of the free parameters in the profile can easily mislead the prediction of the inner density slope.
V Discussion
@@
When we discuss cores, we use the widely adopted definition of density profiles that are asymptotically flat, i.e., whose inner slope $\mathrm{d}\rho/\mathrm{d}r\rightarrow 0$ as $r\rightarrow 0$. That is, we consider a density profile such as the Plummer profile (Equation 3), or the double-power-law profile (Equation 7) with $\gamma=0$, to be cored, even though $\mathrm{d}\rho/\mathrm{d}r<0$ for any finite $r$. Applying the Eddington inversion method to rule out cuspy potentials may require a more limited definition ($\mathrm{d}\rho/\mathrm{d}r>0$ at some finite radius) by which Plummer profiles would not be considered cored. However, we have shown that for typical UFDs, measurements of the true inner slope from small numbers of stars are ambiguous.
@@
A known constraint of the Eddington inversion method (Sánchez Almeida et al., 2024a) is that perfect stellar cores with $\gamma\leq 0$ are incompatible with NFW potentials when the velocity distribution is isotropic or radially biased ($\beta_{\mathrm{ani}}\geq 0$). The incompatibility arises from the method producing a DF that is not everywhere nonnegative (i.e., there exists an $\mathcal{E}$ such that $f(\mathcal{E})<0$) and is hence unphysical.
@@
If we consider our ICs, which are by definition ergodic and spherically symmetric, the DF is solely a function of the relative energy (as in Equation 4). Effectively, $f$ tells us the probability of finding a stellar particle in the system with an energy in the infinitesimal volume $\mathrm{d}\mathcal{E}$ about $\mathcal{E}$. In practice, however, we are unable to numerically sample from such an infinitesimal volume, and rather sample from a coarsened DF $\bar{f}(\mathcal{E})$ (Binney and Tremaine, 2008) that defines a finite volume of width $\Delta\mathcal{E}$, where $\bar{f}$ is the expectation value, or mean, of $f$ within the volume $\Delta\mathcal{E}$. Critically, whilst $\bar{f}(\mathcal{E})$ may be everywhere nonnegative in the volume $\Delta\mathcal{E}$, $f(\mathcal{E})$ in a subregion $\mathrm{d}\mathcal{E}\in\Delta\mathcal{E}$ may not be. The task then becomes ensuring that our approximation
@@
is a valid one over the domain $\Omega$, so that the structure of $\bar{f}$ captures the overall structure of $f$.
@@
In creating our ICs, we calculate $f$ at $2\times 10^{3}$ equally-spaced values of $\mathcal{E}$: this defines our coarsened DF [^1] Note that we use interpolation so as to evaluate our DF at any $\mathcal{E}$, and not just at the values used to construct the DF.. We have tested the effect of increasing the number of $\mathcal{E}$ values $f$ is evaluated at by a factor of 10, and performed a two-sample Kolmogorov–Smirnov (KS) test with a two-tailed null hypothesis to assess if there is a meaningful difference between evaluations of the DF. We find that the test statistic is $t_{\mathrm{KS}}=9.5\times 10^{-4}$ and the $p$-value is $\sim 1.0$: the DFs are indistinguishable. Consequently, we are confident that our constructed $\bar{f}$ is a valid approximation to the true $f$. This is further supported by the constructed ICs agreeing well with the desired Plummer profile (shown in Figure 4), even with no particles exactly in the center requiring high enough energies to result in a negative value being sampled for the DF [^2] Sampling a particle with such high energy is possible but highly unlikely.. While negative DFs are clearly unphysical, analytic DFs are only descriptive, which real and finite systems are not required to fully sample.
Although the combination of perfectly cored density components and perfectly cuspy potentials may lead to partially negative DFs, this does not necessarily lead to real-world contradictions. Any smooth density profile (or DF) is only an approximative description of a finite stellar system, rather than its inherent properties. It is worth pointing out that the “cusp” in an NFW-halo is itself unphysical, but nevertheless, the model provides a good description of the measurable DM distribution. The important question is not whether a model is unphysical at its extremes, but whether it successfully describes the physical reality.
VI Conclusions
@@
The existence of cored rather than cuspy DM profiles in UFDs would have profound implications, effectively falsifying the standard cosmological model. However, our results challenge previous assertions of evidence for DM cores based on cored stellar density profiles. First, we have shown that cored stellar systems can persist within cuspy halos, provided that they have formed in initial equilibrium, as would be expected for a stellar system forming within an existing DM halo in the standard galaxy formation paradigm.
@@
Second, we have demonstrated the limitations of determining the crucial inner slope of the stellar density profiles of UFDs. Even with a conservative sample size and in the absence of observational errors, it is exceptionally difficult to distinguish in UFDs a mildly positive, negative, or zero inner density gradient. This is heightened by the small collection of stars that constitute such a dwarf galaxy, increasing the observed variance in stellar mass density profiles. While future observations may yet lead to a viable application of the Eddington inversion method, we conclude that based on current observations, the existence of stellar cores in UFDs provides no evidence against cuspy DM halos.
@@
We thank the reviewer for their valuable feedback that helped us improve our manuscript. J.H. acknowledges the support of the Magnus Ehrnrooth Foundation. A.R. acknowledges the support of the University of Helsinki Research Foundation. J.H. and T.S. acknowledge support by Research Council of Finland grants 354905 and 339127. J.H., A.R., and T.S. acknowledge support by ERC Consolidator Grant KETJU (no. 818930). M.G.W. acknowledges support from the National Science Foundation (NSF) grant AST-2206046. This work used facilities hosted by the CSC – IT Centre for Science, Finland. Open access funded by Helsinki University Library.
{contribution}
JH: data curation, formal analysis, investigation, methodology, visualization, writing – original draft. AR: formal analysis, investigation, methodology, software, writing – original draft. TS: conceptualization, supervision, writing – original draft. MGW: conceptualization, writing – original draft.
@@ Appendix A Binning of stellar particles
We test the effect of using different numbers of bins when constructing radial profiles on our conclusions, in particular for the inference of the $(\alpha,\beta,\gamma)$ double-power-law profile. In all cases, we use bins of equal particle counts (as opposed to equal bin widths) to ensure sufficient and comparable sample sizes within each bin from which we determine statistics, such as the mean density within a bin. In the following, we discuss signal-to-noise (S/N) in the Poisson sense, i.e., $\mathrm{S/N}=N/\sqrt{N}=\sqrt{N}$.
We first note that determining an optimal binning strategy is a result of two competing factors (e.g., Tarumi et al., 2021): increasing S/N, achieved with increasing the number of particles per bin, and increasing spatial resolution, achieved with increasing the total number of bins. Motivated by observational campaigns of UFDs (e.g., Drlica-Wagner et al., 2015; Tan et al., 2025), we require that the minimum S/N in each radial bin is $\mathrm{S/N}\simeq 10\implies N\simeq 100$. We thus try binning the inner 95% of stellar particles using 5, 10, 25, and 50 bins of equal particle counts [^3] We also tested a case using 95 bins, however due to the low S/N the profile fits were unable to be recovered.. Already we note that using 5 bins, whilst achieving a high S/N of $\sim 30$, barely resolves the inner regions of the stellar profile $r<r_{\mathrm{S}}$, thus setting a lower limit to the number of bins. Using 50 bins achieves our desired $\mathrm{S/N}\simeq 10$, thus setting our upper limit.
@@
A smaller $d_{\mathrm{M}}$ is indicative of a better recovery of the a priori truth than a higher $d_{\mathrm{M}}$. We find that as a function of bin count, $d_{\mathrm{M}}$ decreases exponentially as $d_{\mathrm{M}}\propto N_{\mathrm{bin}}^{-0.43}$: from $N_{\mathrm{bin}}=25$ to $N_{\mathrm{bin}}=50$, the difference in distance measures is only $\Delta d_{\mathrm{M}}=0.24$, indicating the two strategies recover the a priori truth comparatively well. We thus adopt the binning that provides the higher spatial resolution, namely $N_{\mathrm{bin}}=50$ with 95 particles in each bin, as the chosen binning strategy in this work.
@@ Appendix B Fitting density profiles
We fit the $(\alpha,\beta,\gamma)$ double-power-law profile using hierarchical Bayesian inference implemented in Stan (Stan Development Team, 2018), with four chains of 2000 posterior draws each (i.e., post-burn-in period). Recognizing that the 10 fiducial simulations describe the same physical system, albeit with different realizations of the particle sampling, we can jointly model the individual density profile parameters, denoted $\bm{\theta}\equiv\left[\log_{10}\rho_{\mathrm{S}},\log_{10}r_{\mathrm{S}},\log_{10}\alpha,\beta,\gamma\right]$, as realizations of a population of possible density profiles that satisfy the exchangeability criterion (Gelman et al., 2015). The population distribution of density profiles is governed by hyperparameters on each of the latent (or observed) parameters in $\bm{\theta}$ that themselves have weakly informative priors. The prior distributions for the hyperparameters are given in Table 1, and a graphical representation of the Bayesian model in the form of a directed acyclic graph is given in Figure 5. To each latent parameter in $\bm{\theta}$ we assign a normal distribution, with mean $\mu_{X}$ and standard deviation $\sigma_{X}$ for the respective variable $X\in\bm{\theta}$. To capture any covariance between latent parameters, we directly model the correlation matrix $\Omega=L_{\mathrm{c}}L_{\mathrm{c}}^{\top}$, where $L_{\mathrm{c}}$ is the Cholesky factor [^4] The Cholesky factor of a positive-definite matrix is a lower-triangular matrix $L$ such that $\Sigma=LL^{\top}$. of the correlation matrix. We assign a Lewandowski et al. (2009) prior on $L_{\mathrm{c}}$ as $L_{\mathrm{c}}\sim\mathrm{LKJCholesky}(\eta)$, with the prior on the correlation matrix thus becoming $p(\Omega|\eta)\propto\det(\Omega)^{\eta-1}$. The variable $\eta$ controls the concentration of the prior about the identity matrix: by setting $\eta=2$ (Table 1), we mildly favor weak correlations while allowing moderate correlation structure.
@@
After the posterior sampling, we perform standard convergence and posterior-predictive checks to ensure the fits are robust by enforcing a diagnostic value $\hat{R}$ (comparison of between-chain and within-chain variance) of less than 1.05, a high effective sample, and that the number of divergent transitions is less than 1% (for details, see Vehtari et al., 2021; Stan Development Team, 2018). We perform an additional prior sensitivity analysis using the power-scaling method described in Kallioinen et al. (2024), ensuring that the sensitivity diagnostic is less than 0.05 for the latent parameter $\bm{\theta}$ of the model. In doing so, we quantitatively ensure that the obtained posterior distributions are not strongly dependent on the assumed prior distributions.
As our model takes a number of free parameters, degeneracies between parameters are possible, thus highlighting the importance of using a Bayesian method to propagate parameter uncertainty. Typically, we find that the inner slope $\gamma$ is strongly negatively correlated with the characteristic density $\log_{10}\rho_{\mathrm{S}}$, as well as being correlated to a lesser degree with both $\alpha$ (positively) and $\beta$ (negatively). Consequently, $\log_{10}\rho_{\mathrm{S}}$ tends to be negatively correlated with $\alpha$ and positively correlated with $\beta$.